# Beer Path Problems in Temporal Graphs

### v1

@@ 4 Beer path queries with preprocessing
We assume that each query contains as input a source node $x$, a target node $y$, a time interval $[t_{\alpha},t_{\omega}]$ and a Boolean array $A=[a_{1},\ldots,a_{k}]$, with $a_{i}=\texttt{true}$ iff beer vertex $b_{i}$ is active. For simplicity, we assume that each beer vertex $b\in B$ is either active or inactive for the entire duration of a query. We also assume that the set $B$ containing the $k$ beer vertices is known during preprocessing; these vertices may be active or inactive for different queries, but no vertex $v\notin B$ may become a beer vertex.
@@ 4.2 Preprocessing through graph transformation
In this subsection we adopt the graph transformation approach proposed in [20] as preprocessing. This technique converts a temporal graph $G=(V,E)$ into a static, weighted, directed acyclic graph $\tilde{G}=(\tilde{V},\tilde{E})$ that explicitly encodes temporal constraints. The key idea is to model the timing of events in the temporal graph by creating time-stamped copies of each node and connecting them with appropriately weighted edges. The transformed graph $\tilde{G}$ contains $\mathcal{O}(M)$ vertices and edges and can be produced from $G$ in $\mathcal{O}(M\log(d_{max}))$ time. For each $v\in V$, the transformation creates two sets of time-stamped nodes: $\tilde{V}_{in}(v)$, which includes a node $(v,t)$ for every unique time $t$ at which an edge arrives at $v$, and $\tilde{V}_{out}(v)$, which includes a node $(v,t)$ for every time $t$ at which an edge departs from $v$. These sets are sorted in descending time order. Figures 2(a),(b) show a temporal graph $G$ and its transformed graph $\tilde{G}$, respectively. We defer the reader to Appendix C for more details regarding this transformation.
@@ 4.2.1 Fastest beer path on transformed graph
For each vertex $(v,t)\in\tilde{G}$, we keep track of a $3$-valued flag, indicating whether $(v,t)$ has been visited by a beer path (value $2$) or a regular path (value $1$) or not at all (value 00), with higher values superseding smaller ones. We run BFSs from each $(x,t)\in S$ in descending order of $t$ (as described above), setting the flag of $(x,t)$ to $1$. We only visit a vertex if its flag is smaller than the parent’s flag; in this case, the child obtains the parent’s flag. If the child is a copy of some $b_{i}\in B$ with $a_{i}=\texttt{true}$, we always update its flag to $2$. When the flag of a vertex becomes $2$, we calculate the duration of the respective beer path as the difference between the timestamp of that vertex and the timestamp of the source.
@@ Appendix A Omitted Algorithms
1:A temporal graph $G=(V,E)$ in edge stream representation, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$.
2:For each $v\in V$, the list $L_{v}$ of non-dominated $(s,a)$ pairs (from $x$ to $v$) within $[t_{\alpha},t_{\omega}]$.
@@
1:A temporal graph $G=(V,E)$ in edge stream representation, source vertex $x$, time interval $[t_{\alpha},t_{\omega}]$.
2:For each $v\in V$, the list $S_{v}$ of distance-wise non-dominated $(d,a)$ pairs (from $x$ to $v$) within $[t_{\alpha},t_{\omega}]$.
@@ A.2 Modifed non-dominated path algorithm for new domination criterion
1:A temporal graph $G=(V,E)$ in reverse edge stream representation, target vertex $y$, time interval $[t_{\alpha},t_{\omega}]$.
2:For each $v\in V$, the list $I_{v}$ of inverse-distance-wise non-dominated $(d,s)$ pairs (from $v$ to $y$) within $[t_{\alpha},t_{\omega}]$.
@@ A.3 Omitted MTLDP and latest-departure beer path algorithms
Algorithm 9 MTLDP$(G,[t_{\alpha},t_{\omega}],fin)$
1:A temporal graph $G$ in its reverse edge stream representation, a time window $[t_{\alpha},t_{\omega}]$ and final times $fin[v]$, such that $t_{\alpha}\leq fin[v]\leq t_{\omega}$ or $fin[v]=-\infty$, for all $v\in V$.
2:The MTLDP times within $[t_{\alpha},t_{\omega}]$ to every vertex, assuming we can arrive at time $fin[v]$ at each $v\in V$.
3:$\tau[v]\leftarrow fin[v]$ for all $v\in V$
@@
Algorithm 10 MTLDP_alternative$(G,[t_{\alpha},t_{\omega}],fin)$
1:A temporal graph $G$ in adjacency list representation, a time window $[t_{\alpha},t_{\omega}]$ and final times $fin[v]$, such that $t_{\alpha}\leq fin[v]\leq t_{\omega}$ or $fin[v]=-\infty$, for all $v\in V$.
2:The MTLDP times within $[t_{\alpha},t_{\omega}]$ to every vertex, assuming we can arrive at time $fin[v]$ at each $v\in V$.
3:$\tau[v]\leftarrow fin[v]$ for all $v\in V$
4:$PQ\leftarrow\emptyset$triangleright a max-priority queue
5:$PQ.enqueue(\tau[v],v)$ for all $v\in V$triangleright Add $v$ to priority queue with $\tau[v]$ priority
@@
9:         Let $(u,v,t,\lambda)\in E(u,v)$ be a temporal edge with maximum $t$ s.t. $t+\lambda\leq\tau[v]$triangleright binary search on $t+\lambda$ of $e\in E(u,v)$
@@
Algorithm 11 LDBP$(G,x,[t_{\alpha},t_{\omega}])$
1:A temporal graph $G$, source vertex $x$, a time window $[t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
2:The LDBP time to $y$ from every vertex $v\in V$ within $[t_{\alpha},t_{\omega}]$.
3:$fin[y]\leftarrow t_{\omega},\,init[v]\leftarrow-\infty$ for all $v\in V\setminus\{x\}$
4:$\tau[v]\leftarrow\texttt{MTLDP($G,[t_{\alpha},t_{\omega}],fin$)}$, for $v\in V$
5:$fin^{\prime}[v]\leftarrow-\infty$ for all $v\in V$
@@
7:     $beer\_time\leftarrow\max\{time\in T_{b}\mid time\leq\tau[b]\}$triangleright $\infty$- ∞ if it does not exist
@@
10:$\tau^{\prime}[v]\leftarrow\texttt{MTLDP($G,[t_{\alpha},t_{\omega}],fin^{\prime}$)}$, for $v\in V$
@@ Appendix B Omitted proofs
We now prove the correctness of Algorithm 5 by induction. We only describe the induction briefly, since it is similar to the proof of correctness of Dijkstra’s classic algorithm. For each $v\in V$ let $\tau[v]$ be the time calculated by the algorithm and $\tau^{*}[v]$ be the respective MSEAP time within $[t_{\alpha},t_{\omega}]$. By definition of MSEAP, for the node with minimum $init[u]$ it is $\tau^{*}[u]=init[u]$; thus, the first node extracted from the queue has $\tau[u]=\tau^{*}[u]$. For each $v\in N_{out}(u)$, $\tau[v]$ is then updated through the $u-v$ edge that arrives at $v$ at the earliest possible time. Assuming the graph has no dominated edges, this is an edge with minimum $t$ s.t. $t\geq\tau[u]$. Inductively, any $u\in V$ extracted from the queue has $\tau[u]=\tau^{*}[u]$. This happens because paths that only use processed nodes have been considered in previous steps and it is impossible to arrive at $u$ earlier than $\tau[u]$ by traversing an unprocessed node $v\in V$ with $\tau[v]\geq\tau[u]$. Regarding the update of $\tau[v]$ for $v\in N_{out}(u)$, the same things hold as in the base of the induction, thus completing the proof.
@@
If there is a temporal $x-y$ beer path within time interval $[t_{\alpha},t_{\omega}]$, then there is an $x-y$ FBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, such that its subpaths $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ are non-dominated paths.
@@
Let there be a temporal beer path $P=\langle u_{0}=x,u_{1},\ldots,u_{j^{\prime}}=b,\ldots,u_{p^{\prime}}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$. Define the subpaths of $P$: $Q=\langle u_{0}=x,u_{1},\ldots,u_{j^{\prime}}=b\rangle$, $R=\langle u_{j^{\prime}}=b,u_{j^{\prime}+1},\ldots,u_{p^{\prime}}=y\rangle$. If $Q$ is a dominated path, then there is some other path $Q^{\prime}=\langle w_{0}=x,w_{1},\ldots,w_{f}=b\rangle$ such that $start(Q^{\prime})\geq start(Q)$ and $end(Q^{\prime})\leq end(Q)$, with at least one of the inequalities being strict. Concatenate $Q^{\prime}$ with $R$ to obtain a $x-y$ path $P^{\prime}$. By the first inequality we obtain $dura(P^{\prime})\leq dura(P)$. By the second inequality we infer that $P^{\prime}$ also traverses $b$ at time $t$ and that $P^{\prime}$ is a valid temporal path. Thus, $P^{\prime}$ is a beer path with smaller or equal duration than $P$ and is also within $[t_{\alpha},t_{\omega}]$. Applying a similar argument for $R$ completes the proof.
@@
If there is a temporal $x-y$ beer path within time interval $[t_{\alpha},t_{\omega}]$, then there is a $x-y$ SBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, s.t. $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ is a distance-wise non-dominated path (within $[t_{\alpha},t_{\omega}]$) and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ is an inverse-distance-wise non-dominated path (within $[t_{\alpha},t_{\omega}]$).
@@
Let there be a temporal beer path $P=\langle u_{0}=x,u_{1},\dots,u_{j^{\prime}}=b,\dots,u_{p^{\prime}}=y\rangle$ within $[t_{\alpha},t_{\omega}]$, traversing $b\in B$ at some time $t\in T_{b}$. Define the sub-paths of $P$: $Q=\langle u_{0}=x,u_{1},\dots,u_{j^{\prime}}=b\rangle$ and $R=\langle u_{j^{\prime}}=b,u_{j^{\prime}+1},\dots,u_{p^{\prime}}=y\rangle$. If $Q$ is distance-wise dominated (within $[t_{\alpha},t_{\omega}]$), there exists a path $Q^{\prime}=\langle w_{0}=x,w_{1},\dots,w_{f}=b\rangle$ (within $[t_{\alpha},t_{\omega}]$), such that $\text{dist}(Q^{\prime})\leq\text{dist}(Q)$ and $\text{end}(Q^{\prime})\leq\text{end}(Q)$, with at least one of the two inequalities strict. Concatenate $Q^{\prime}$ with $R$ to obtain a path $P^{\prime}$ from $x$ to $y$. Because of the first inequality we get $\text{dist}(P^{\prime})\leq\text{dist}(P)$. By the second inequality we infer that $P^{\prime}$ also traverses $b$ at time $t$ and is a valid temporal path. Hence, $P^{\prime}$ is a beer path within $[t_{\alpha},t_{\omega}]$ whose total distance is smaller than or equal to that of $P$.
If $R$ is inverse-distance-wise dominated (within $[t_{\alpha},t_{\omega}]$), there is a path $R^{\prime}=\langle z_{0}=b,z_{1},\dots,z_{g}=y\rangle$ (within $[t_{\alpha},t_{\omega}]$) such that $\text{dist}(R^{\prime})\leq\text{dist}(R)$ and $\text{start}(R^{\prime})\geq\text{start}(R)$, again with at least one inequality being strict. Concatenate $Q^{\prime}$ with $R^{\prime}$ to obtain a $x-y$ path $P^{\prime\prime}$. By the first inequality we obtain $\text{dist}(P^{\prime\prime})\leq\text{dist}(P^{\prime})$. By the second inequality, we obtain that $P^{\prime\prime}$ is a valid temporal beer path. Therefore, $P^{\prime\prime}$ is a beer path within $[t_{\alpha},t_{\omega}]$ whose total distance is smaller than or equal to that of $P^{\prime}$.
@@
Algorithm 7 computes all inverse-distance-wise non-dominated $(d,s)$ pairs from each $v\in V$ to $y$ within $[t_{\alpha},t_{\omega}]$, in time $\mathcal{O}(n+M\log(d_{out}))$.
@@
Let $P=\langle v_{0}=v,v_{1},\dots,v_{p}=y\rangle$ be an inverse-distance-wise non-dominated path from $v$ to $y$ within $[t_{\alpha},t_{\omega}]$. For $i\in\{0,\ldots,p-1\}$, let $s_{i}$ be the time in which $P$ departs from $v_{i}$ and $d_{i}$ be the distance between $v_{i}$ and $y$ in $P$. We will prove that $(d_{0},s_{0})$ is contained in the list $I_{v}$ returned by Algorithm 7. Since we use a reverse edge stream, the edges of $P$ are processed in reverse order (starting from $y$). We use induction on the edges of $P$.
@@
From the above induction we infer that $I_{v}$ contains an element $(d,s)$ with $d\leq d_{0}$ and $s\geq s_{0}$. By assumption, $P$ is an inverse-distance-wise non-dominated path within $[t_{\alpha},t_{\omega}]$, which implies that neither of the above inequalities can be strict. Hence, $I_{v}$ contains $(d_{0},s_{0})$, which finishes the correctness proof.
@@
In this section we describe in detail the graph transformation of [20], which converts a temporal graph $G=(V,E)$ into an equivalent static, weighted, directed acyclic graph $\tilde{G}=(\tilde{V},\tilde{E})$.

### v2

@@ 4 Beer path queries with preprocessing
We assume that each query contains as input a source node $x$, a target node $y$, a time interval $t_{\alpha},t_{\omega}]$ and a Boolean array $A=[a_{1},\ldots,a_{k}]$, with $a_{i}=\texttt{true}$ iff beer vertex $b_{i}$ is active. For simplicity, we assume that each beer vertex $b\in B$ is either active or inactive for the entire duration of a query. We also assume that the set $B$ containing the $k$ beer vertices is known during preprocessing; these vertices may be active or inactive for different queries, but no vertex $v\notin B$ may become a beer vertex.
@@ 4.2 Preprocessing through graph transformation
[…]
@@ 4.2.1 Fastest beer path on transformed graph
For each vertex $(v,t)\in\tilde{G}$, we keep track of a $3$-valued flag, indicating whether $(v,t)$ has been visited by a beer path (value $2$) or a regular path (value $1$) or not at all (value 0), with higher values superseding smaller ones. We run BFSs from each $(x,t)\in S$ in descending order of $t$ (as described above), setting the flag of $(x,t)$ to $1$. We only visit a vertex if its flag is smaller than the parent’s flag; in this case, the child obtains the parent’s flag. If the child is a copy of some $b_{i}\in B$ with $a_{i}=\texttt{true}$, we always update its flag to $2$. When the flag of a vertex becomes $2$, we calculate the duration of the respective beer path as the difference between the timestamp of that vertex and the timestamp of the source.
@@ Appendix A Omitted Algorithms
1:A temporal graph $G=(V,E)$ in edge stream representation, source vertex $x$, time interval $t_{\alpha},t_{\omega}]$.
2:For each $v\in V$, the list $L_{v}$ of non-dominated $(s,a)$ pairs (from $x$ to $v$) within $t_{\alpha},t_{\omega}]$.
@@
1:A temporal graph $G=(V,E)$ in edge stream representation, source vertex $x$, time interval $t_{\alpha},t_{\omega}]$.
2:For each $v\in V$, the list $S_{v}$ of distance-wise non-dominated $(d,a)$ pairs (from $x$ to $v$) within $t_{\alpha},t_{\omega}]$.
@@ A.2 Modifed non-dominated path algorithm for new domination criterion
1:A temporal graph $G=(V,E)$ in reverse edge stream representation, target vertex $y$, time interval $t_{\alpha},t_{\omega}]$.
2:For each $v\in V$, the list $I_{v}$ of inverse-distance-wise non-dominated $(d,s)$ pairs (from $v$ to $y$) within $t_{\alpha},t_{\omega}]$.
@@ A.3 Omitted MTLDP and latest-departure beer path algorithms
[…]
1:A temporal graph $G$ in its reverse edge stream representation, a time window $t_{\alpha},t_{\omega}]$ and final times $fin[v]$, such that $t_{\alpha}\leq fin[v]\leq t_{\omega}$ or $fin[v]=$-$\infty$, for all $v\in V$.
2:The MTLDP times within $t_{\alpha},t_{\omega}]$ to every vertex, assuming we can arrive at time $fin[v]$ at each $v\in V$.
[…]
@@
[…]
1:A temporal graph $G$ in adjacency list representation, a time window $t_{\alpha},t_{\omega}]$ and final times $fin[v]$, such that $t_{\alpha}\leq fin[v]\leq t_{\omega}$ or $fin[v]=$-$\infty$, for all $v\in V$.
2:The MTLDP times within $t_{\alpha},t_{\omega}]$ to every vertex, assuming we can arrive at time $fin[v]$ at each $v\in V$.
[…]
@@
9:   Let $(u,v,t,\lambda)\in E(u,v)$ be a temporal edge with maximum $t$ s.t. $t+\lambda\leq\tau[v]$ $\triangleright$ binary search on $t+\lambda$ of $e\in E(u,v)$
@@
[…]
1:A temporal graph $G$, source vertex $x$, a time window $t_{\alpha},t_{\omega}]$, a set of beer vertices $B$ and a list $T_{i}$ of time units for each of them.
2:The LDBP time to $y$ from every vertex $v\in V$ within $t_{\alpha},t_{\omega}]$.
[…]
@@
7:  $beer\_time\leftarrow\max\{time\in T_{b}\mid time\leq\tau[b]\}$ $\triangleright$ -$\infty$ if it does not exist
@@
[…]
@@ Appendix B Omitted proofs
We now prove the correctness of Algorithm 5 by induction. We only describe the induction briefly, since it is similar to the proof of correctness of Dijkstra’s classic algorithm. For each $v\in V$ let $\tau[v]$ be the time calculated by the algorithm and $\tau^{*}[v]$ be the respective MSEAP time within $t_{\alpha},t_{\omega}]$. By definition of MSEAP, for the node with minimum $init[u]$ it is $\tau^{*}[u]=init[u]$; thus, the first node extracted from the queue has $\tau[u]=\tau^{*}[u]$. For each $v\in N_{out}(u)$, $\tau[v]$ is then updated through the uvu-v edge that arrives at $v$ at the earliest possible time. Assuming the graph has no dominated edges, this is an edge with minimum $t$ s.t. $t\geq\tau[u]$. Inductively, any $u\in V$ extracted from the queue has $\tau[u]=\tau^{*}[u]$. This happens because paths that only use processed nodes have been considered in previous steps and it is impossible to arrive at $u$ earlier than $\tau[u]$ by traversing an unprocessed node $v\in V$ with $\tau[v]\geq\tau[u]$. Regarding the update of $\tau[v]$ for $v\in N_{out}(u)$, the same things hold as in the base of the induction, thus completing the proof.
@@
If there is a temporal xyx-y beer path within time interval $t_{\alpha},t_{\omega}]$, then there is an xyx-y FBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, such that its subpaths $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ are non-dominated paths.
@@
Let there be a temporal beer path $P=\langle u_{0}=x,u_{1},\ldots,u_{j^{\prime}}=b,\ldots,u_{p^{\prime}}=y\rangle$ within $t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$. Define the subpaths of $P$: $Q=\langle u_{0}=x,u_{1},\ldots,u_{j^{\prime}}=b\rangle$, $R=\langle u_{j^{\prime}}=b,u_{j^{\prime}+1},\ldots,u_{p^{\prime}}=y\rangle$. If $Q$ is a dominated path, then there is some other path $Q^{\prime}=\langle w_{0}=x,w_{1},\ldots,w_{f}=b\rangle$ such that $start(Q^{\prime})\geq start(Q)$ and $end(Q^{\prime})\leq end(Q)$, with at least one of the inequalities being strict. Concatenate $Q^{\prime}$ with $R$ to obtain a xyx-y path $P^{\prime}$. By the first inequality we obtain $dura(P^{\prime})\leq dura(P)$. By the second inequality we infer that $P^{\prime}$ also traverses $b$ at time $t$ and that $P^{\prime}$ is a valid temporal path. Thus, $P^{\prime}$ is a beer path with smaller or equal duration than $P$ and is also within $t_{\alpha},t_{\omega}]$. Applying a similar argument for $R$ completes the proof.
@@
If there is a temporal xyx-y beer path within time interval $t_{\alpha},t_{\omega}]$, then there is a xyx-y SBP $\langle v_{0}=x,v_{1},\ldots,v_{j}=b,\ldots,v_{p}=y\rangle$ within $t_{\alpha},t_{\omega}]$, traversing $b\in B$ at a time $t\in T_{b}$, s.t. $\langle v_{0}=x,v_{1},\ldots,v_{j}=b\rangle$ is a distance-wise non-dominated path (within $t_{\alpha},t_{\omega}]$) and $\langle v_{j}=b,v_{j+1},\ldots,v_{p}=y\rangle$ is an inverse-distance-wise non-dominated path (within $t_{\alpha},t_{\omega}]$).
@@
Let there be a temporal beer path $P=\langle u_{0}=x,u_{1},\dots,u_{j^{\prime}}=b,\dots,u_{p^{\prime}}=y\rangle$ within $t_{\alpha},t_{\omega}]$, traversing $b\in B$ at some time $t\in T_{b}$. Define the sub-paths of $P$: $Q=\langle u_{0}=x,u_{1},\dots,u_{j^{\prime}}=b\rangle$ and $R=\langle u_{j^{\prime}}=b,u_{j^{\prime}+1},\dots,u_{p^{\prime}}=y\rangle$. If $Q$ is distance-wise dominated (within $t_{\alpha},t_{\omega}]$), there exists a path $Q^{\prime}=\langle w_{0}=x,w_{1},\dots,w_{f}=b\rangle$ (within $t_{\alpha},t_{\omega}]$), such that $\text{dist}(Q^{\prime})\leq\text{dist}(Q)$ and $\text{end}(Q^{\prime})\leq\text{end}(Q)$, with at least one of the two inequalities strict. Concatenate $Q^{\prime}$ with $R$ to obtain a path $P^{\prime}$ from $x$ to $y$. Because of the first inequality we get $\text{dist}(P^{\prime})\leq\text{dist}(P)$. By the second inequality we infer that $P^{\prime}$ also traverses $b$ at time $t$ and is a valid temporal path. Hence, $P^{\prime}$ is a beer path within $t_{\alpha},t_{\omega}]$ whose total distance is smaller than or equal to that of $P$.
If $R$ is inverse-distance-wise dominated (within $t_{\alpha},t_{\omega}]$), there is a path $R^{\prime}=\langle z_{0}=b,z_{1},\dots,z_{g}=y\rangle$ (within $t_{\alpha},t_{\omega}]$) such that $\text{dist}(R^{\prime})\leq\text{dist}(R)$ and $\text{start}(R^{\prime})\geq\text{start}(R)$, again with at least one inequality being strict. Concatenate $Q^{\prime}$ with $R^{\prime}$ to obtain a xyx-y path $P^{\prime\prime}$. By the first inequality we obtain $\text{dist}(P^{\prime\prime})\leq\text{dist}(P^{\prime})$. By the second inequality, we obtain that $P^{\prime\prime}$ is a valid temporal beer path. Therefore, $P^{\prime\prime}$ is a beer path within $t_{\alpha},t_{\omega}]$ whose total distance is smaller than or equal to that of $P^{\prime}$.
@@
Algorithm 7 computes all inverse-distance-wise non-dominated $(d,s)$ pairs from each $v\in V$ to $y$ within $t_{\alpha},t_{\omega}]$, in time $\mathcal{O}(n+M\log(d_{out}))$.
@@
Let $P=\langle v_{0}=v,v_{1},\dots,v_{p}=y\rangle$ be an inverse-distance-wise non-dominated path from $v$ to $y$ within $t_{\alpha},t_{\omega}]$. For $i\in\{0,\ldots,p-1\}$, let $s_{i}$ be the time in which $P$ departs from $v_{i}$ and $d_{i}$ be the distance between $v_{i}$ and $y$ in $P$. We will prove that $(d_{0},s_{0})$ is contained in the list $I_{v}$ returned by Algorithm 7. Since we use a reverse edge stream, the edges of $P$ are processed in reverse order (starting from $y$). We use induction on the edges of $P$.
@@
From the above induction we infer that $I_{v}$ contains an element $(d,s)$ with $d\leq d_{0}$ and $s\geq s_{0}$. By assumption, $P$ is an inverse-distance-wise non-dominated path within $t_{\alpha},t_{\omega}]$, which implies that neither of the above inequalities can be strict. Hence, $I_{v}$ contains $(d_{0},s_{0})$, which finishes the correctness proof.
@@
[…]